OpenStudy (anonymous):

Simplify: 2s^2-5s-12 ----------- 2s^2-9s+4 Please show your work so I can understand what you're doing.

OpenStudy (compassionate):

Hi, my name is Stephen and I'll be your tutor. The first step to solving algebraic fractions is simplifying. There is nothing ti simplify in this expression so we will move onto the second step. In order to reduce the fraction we must subtract the common terms from the bottom to the top. The general rule is: \[\frac{ x^3 + y^3 }{ x + y } = \frac{ x^2 + y^2}{ 1 }\] You can see we took the x and y from the bottom, and subtracted it from the top x-cubed and y-cubed terms. NOTE: You can ONLY subtract common terms. \[\frac{ 2s^2 - 5s - 12 }{ 2s^2 - 9s + 4 }\] In the above equation you want to subtract your squard terms, your single terms (-5s and -9s) and your numbers (-12 and +4)

OpenStudy (anonymous):

I don't really understand what you're doing because my teacher told me that to simplify it we need to factor it, then we could subtracts common terms. Thats how I've been learning it.

OpenStudy (anonymous):

hi yes you do it similar to what you learned in class, you get the factors of each expression and then 'cancel' what is common on the top and the bottom as shown here:\[\frac{ (2s+3)(s-4) }{ (2s-1)(s-4) }=\frac{ 2s+3 }{ 2s-1}\]

OpenStudy (anonymous):

Thank you!(: